OpenStudy (anonymous):

Write a system of two equations using x and y for the following problem: The sum of two numbers is 170 and their difference is 26. Find the two numbers.

OpenStudy (anonymous):

Is this right? x+y=170 x-y=26

OpenStudy (anonymous):

The 2 numbers could be expressed as 'x' and 'y' so you can put x + y = 170 and also if they have a difference of 26 you can write x - y = 26.

OpenStudy (anonymous):

Yay I was right!

OpenStudy (anonymous):

How do I find the two numbers?

OpenStudy (anonymous):

Write a system of two equations using x and y in each for the following. The sum of two numbers is 170 and their difference is 26. A + B = 170 A -- B = 26 Answer 98, 72

OpenStudy (anonymous):

You already found the equations, add them together...